7.3 Use ratios to computo the missing data in the following financ al statements X Ltd. Statement of comprehensive income for year to 31/12/2016 Shs Sales Less Gross pro fit Less operating expenses Interest expense 9 20,000

media%2F298%2F29857d04-7211-4f32-ad06-bd

media%2F6e4%2F6e463353-84b5-4278-87a0-e1
Answer #1
Income Statement
Particulars Shs
Sales(Note 7) 100000
Less: Cost of Goods Sold(Note 7) 40000
Gross Profit 60000
Less: Operating expenses 20000
Interest expense(Note 2) 2000
Income before tax(Note 1) 38000
Less: Income tax(50%) 19000
Net income(Note 1) 19000
Balance sheet
Particulars Shs
Assets
Current Assets
Cash(Note 9) 8,000
A/Receivables(Note 8) 10,000
Inventory(Note 6) 18,000
Total Current Assets(Note 5) 36,000
Plant and equipment
Plant and equipment(Note 10) 59000
Total Assets 95,000
Liabilities and stockholders equity
Current liabilities(Note 4) 18000
10% Long term liabilties 20000
Common stock @10/- par value(Note 3) 50000
Retained earnings(Note 4) 7000
Total liability and stockholders equity 95000

Notes:

1. Income before tax=19,000 and tax rate is 50%.

So, income before tax=19,000/0.5=38,000

So, net income=38,000-19,000=19,000

2. Interest expense attributable only to long term liabilities whose value is 20,000 as per balance sheet carrying a rate of 10%.

So, gross profit=profit before tax+interest expense+operating expense

=38,000+2,000+22,000

=60,000

So, interest expense=20,000*10%=2,000

3. Earnings per share=3.8 and so additional capital was issued.

So, number of shares=19,000/3.8=5,000.

Thus, common stock @10/- par value=5,000*10=50,000

4. Debts to total assets ratio=40%.

Now, total liability and stockholders equity=95,000 and in balance sheet total assets will be equal to total liability and stockholders equity.

So, total debts=95,000*40%=38,000 out of which 20,000 is 10% Long term liabilties.

So, current liabilties=38,000-20,000=18,000

Thus, balancing figure is retained earnings.

Retained earnings=95,000-18,000-20,000-50,000=7,000

5. Current ratio is 2 which means that current assets will be twice of current liabilties which is 18,000.

So, current assets=18,000*2=36,000

6. Quick ratio=1 which is (Current Assets-inventory)/Current Liabilities

So, Current Assets-Inventory=Current Liabilities

36,000-Inventory=18,000

Inventory=18,000

7. Inventory turnover=2 times which is Cost of Goods Sold to Average Inventory Ratio.

Now, opening inventory as on 1/1/2016 is 22,000.

Closing inventory as calculated in Note 6 is 18,000.

So, average inventory =(22,000+18,000)/2=20,000

Thus, Cost of Goods Sold=20,000*2=40,000.

So, Sales=Gross Profit+Cost of Goods Sold

=60,000+40,000

=1,00,000

8. Receivables turnover=10 times which is Sales to Average Receivables Ratio.

   Now, opening receivables as on 1/1/2016 is 10,000.

Average receivables=(10,000+Closing receivables)/2

Sales=1,00,000

Thus, 1,00,000/(10,000+Closing receivables)/2=10

Or, 200,000/(10,000+Closing receivables)=10

Or, 10,000+Closing receivables=20,000

Thus, Closing receivables=20,000-10,000=10,000

9. Cash=Current Assets-Inventory-Receivables

=36,000-18,000-10,000

=8,000

10. Plant and equipment is balancing figure in the Balance Sheet.

Total assets=95,000 outr of which current is 36,000.

So, Plant and equipment=95,000-36,000=59,000